OpenStudy (anonymous):

What is the simplified form of the expression x^8 * 2y^10 * 5x^5

OpenStudy (anonymous):

you have two x like terms, multiply the x^8 by the 5x^5 to combine them into one term your y term is as simple as it will get

OpenStudy (anonymous):

I got the answer of 10xy^23. Is that correct?

OpenStudy (anonymous):

no you cannot combine the exponents of different variables

OpenStudy (anonymous):

x^8 * 5x^5 = 5x^13 5x^13 * 2y^10 = 10x^13y^10

OpenStudy (anonymous):

so 2x^2*2y^2*2z^2 would = 8x^2y^2z^2

OpenStudy (anonymous):

You can only combine the coefficients on different variables

OpenStudy (anonymous):

Ok, so 8x^13 y^10?

OpenStudy (anonymous):

5x^13 * 2y^10 = 10x^13y^10 10x^13y^10

OpenStudy (anonymous):

Got it! Thanks!

OpenStudy (vjordillo):

\[10x^{13}y ^{10}\]